Question 4:

Let X be a random variable with the following probability distribution:


X -3 6 9

P(X=x) 1/6 1/2 1/3


If the function of random variable is defined as f(X) = (2X+1)2, find µg(X).



1
Expert's answer
2021-08-09T15:39:15-0400


To find "\mu_{g(X)}" where "g(x) = (2x+1)^2"

Here "\mu_{g(X)}" means expected value of g(x)

"\mu_{g(X)} = \sum g(X) f(X) \\

f(x) = (2x+1)^2 \\

f(-3) = (2(-3) +1)^2 = (-5)^2 = 25 \\

f(6) = (2 \times 6 +1)^2 = 13^2 = 169 \\

f(9) = (2 \times 9 +1)^2 = 19^2= 361 \\

\mu_{g(X)} = 25 \times \frac{1}{6} + 169 \times \frac{1}{2} + 361 \times \frac{1}{3} \\

= 4.16+84.5+120.33 \\

= 208.99 ≈209"